OpenStudy (anonymous):

If I have 2^(3-x) + 2^x = 9 how do I isolate the x's?

OpenStudy (anonymous):

Bring down logs and use logx^n = nlogx

OpenStudy (haseeb96):

2^3.2^-x + 2^x =9 \[\frac{ 2^3 }{ 2^x }+2^x = 9\] \[\frac{ 8+4x^2 }{ 2^x }=9\] 8+4x^2=18^x 4x^2-18^x+8=0 taking two common from this equation 2(2x^2-9x+4)=0 2x^2-9x+4=0 using factorized method 2x^2-1x-8x+4=0 x(2x-1)-4(2x-1)=0 (x-4)(2x-1)=0 either x-4=0 or 2x-1=0 x=4 or 2x=1 x=1\2 this is the correct answer

OpenStudy (anonymous):

Thanks for the help I actually think the answer is 0 and 3 though

OpenStudy (haseeb96):

no, it is wrong

OpenStudy (anonymous):

no its right, 9 * 2^x is not 18^x otherwise your math was right. I saw what I did wrong and it is 0 and 3. You can check it if you plug it back into the equation.